OpenStudy (anonymous):

Help with proving W = (1,2,4x+3Y) not a vector

OpenStudy (anonymous):

are there any more details to this question?

OpenStudy (anonymous):

x and y are real numbers and V = R3

OpenStudy (anonymous):

Help with proving W = (1x,2y,4x+3Y) not a vector , x and y are real numbers and V = R3

OpenStudy (anonymous):

not a vector or not a vector space?

OpenStudy (anonymous):

not a vector sub space

OpenStudy (anonymous):

okay, makes more sense now

OpenStudy (anonymous):

it appears to pass the conditions to be a subspace of R^3

OpenStudy (anonymous):

how to prove with addition

OpenStudy (anonymous):

Let \[\vec{a}=(x_1,2y_2,4x_1+3y_1)\] and \[\vec{b}=(x_2,y_2,4x_2+3y_2)\] be in W

OpenStudy (anonymous):

Now \[\vec{a}+\vec{b}=(x_1+x_2,y_1+y_2,4x_1+3y_1+4x_2+3y_2)\] group and factor the third component to get \[(x_1+x_2,y_1+y_2,4(x_1+x_2)+3(y_1+y_2))\]

OpenStudy (anonymous):

since \[x_1,x_2,y_1,y_2\in \mathbb{R}\] clearly this vector sum of arbitrary vectors in W is in W

OpenStudy (anonymous):

I should have made the statement about \[x_1,x_2,y_1,y_2\in \mathbb{R}\] first of course

OpenStudy (anonymous):

Thank you I am reviewing to make sure I understand.

OpenStudy (anonymous):

would it work the same if W = x,2y,4x - 3Y

OpenStudy (anonymous):

it should, but I have not tried it, Proving a subspace is one of those "turn the crank" type proofs in linear algebra that may seem a little abstract at first but once yo get the process down is straight forward but sometimes tedious...

OpenStudy (anonymous):

Thank You

OpenStudy (anonymous):

IS W subspace of V, W is the set of all functions F(0)=1 , V=C(\[-\infty\],\[\infty\])